# Monadic Second Order Logic of the Real Order with the Set Quantifier ranging over the Borel Sets

### v2

@@ 5. Decidability of MSO over the real order with Borel Set Quantifiers
Recall Notation 4.3. We shall translate $\varphi$ into an MSO-sentence $\varphi^{\ast}$ of ${\mathfrak{M}}_{1+\alpha}$. Let $H_{1},\ldots H_{n}$ be the set variables appearing in $\varphi$. We identify individual variables $\rho$ in $\varphi$ with sets $\{\rho\}$, which are closed sets in $[T]$. In particular they also are $\Sigma_{\alpha}$-sets. Let $\rho_{1},\ldots\rho_{m}$ be all the individual variables appearing in $\varphi$. By applying Lemma 4.8 finitely many times if necessary, we can assume that we have chosen the codes $c(H_{i})$ and $c({\rho}_{j})$ for $i<n,j<m$ within $T^{1+\alpha}]$ so that
@@
Suppose that $\langle F_{\alpha}:\,\alpha<\omega_{1}\rangle$ is a decidable sequence of sentences, namely there is an algorithm which on an input (a code for) $\alpha$ outputs the truth value of $F_{\alpha}$. Then the statement $(\forall\alpha<\omega_{1})F_{\alpha}]$ is decidable.
@@
As the set of decidable sentences is closed under $\neg,\wedge$ and $\vee$, the nontrivial case is the quantifier case. Suppose first that $\varphi$ has the prenex normal form $(\forall{\mathbf{A}})\psi({\mathbf{A}})$ where $\psi$ does not have any other set quantifiers. For each $\alpha<\omega_{1}$ consider the sentence $F_{\alpha}\equiv(\forall{\mathbf{A}})[(\Sigma_{\alpha}({\mathbf{A}})\vee\Pi_{\alpha}({\mathbf{A}})\implies\psi({\mathbf{A}})]$. This is a sentence in the MSO theory of $\mathfrak{C}$ and so decidable by Theorem 5.1.
@@ 6. Effective Borel Determinacy
A winning strategy $S$ for I is a function defined on $\bigcup_{n<\omega}{}^{2n+1}2$ such that for any play $\rho$ of $G(H)$ such that at every step $n$ the player I plays $S(\rho\upharpoonright(2n+2))$, the player I wins (this is a recursive definition). We similarly define a winning strategy for II. The set $H$ is determined if one of the players has a winning strategy in $G(H)$.
@@
Philip Wolfe [21] proved the determinacy of $F_{\sigma}$-sets. In [19, §2.5], Rabin showed that the statement that every $F_{\sigma}$-set is determined, is expressible as an MSO sentence over ${\mathfrak{N}}_{2}$. He concluded that therefore, Wolfe’s theorem is among those decided by the decision procedure for S2S. In particular, knowing that all $F_{\sigma}$-sets are determined, we can conclude that for every $F_{\sigma}$-set there is a recursive procedure to decide which player wins the game. This recursive procedure is given by a run of the Rabin automaton. This is what we may call an effective version of the determinacy for $F_{\sigma}$-sets.
@@ 7. Closing Remarks
The elegant proof of the Borel Determinacy given by Kechris in [9, Th, 20.5] relates to Martin’s paper [15]. Martin mentions in [15] that it was proposed by Yannis Moschovakis in his book [18] that such a proof should exist, although it was not found then. Before this proof, some earlier papers on the subject used a different technique. The idea was to associate to each Borel game $G$ a game $G^{\ast}$ on an open set, so that if the player I or II has a winning strategy in $G^{\ast}$, then the player I or II respectively has a winning strategy in $G$. The Borel Determinacy then follows from the Gale-Stewart theorem on the determinacy of open sets [6]. The idea of translation between $G$ and $G^{*}$ was introduced by Martin in [13], where he claims the proof of Borel Determinacy from the existence of an $\omega_{1}$-Erdős cardinal. His paper [14] claims Borel determinacy in ZFC, using a better coding. In the two papers the details of the proofs are missing or incorrect. With this in mind, we have tried to prove Borel Determinacy in yet another way, to match Martin’s original intuition. This would be by playing games on codes for $\Sigma_{\alpha}\cup\Pi_{\alpha}$-sets, understanding those as closed games in an appropriate Polish subspace of $T^{1+\alpha}]$ and then projecting to the ordinary games on closed sets. The intended proof is by induction on $\alpha$ and uses a direct proof for $F_{\sigma}$-determinacy given in [9, Exercise 20.9] at the non-trivial part of the projection between games in the successor case. This proof would correspond to the original programme of proving Borel Determinacy as intended by Martin in his papers [13] and [14]. We have not worked out all of the details of this proof, as from the purely mathematical point of view it seems unnecessary to give a very long combinatorial argument for a theorem whose proof is already known and well written ([9, Th 20.5]). In fact, we believe that the combinatorial proof is simply a recounting of the decision procedure for the player which has the winning strategy. The part of the combinatorial proof that we did write is available in the code of this paper on the ArXiv. If an interested person uses it, they should ask for a permission and their work should have a proper mention of the source (this paper). The reason we mention these issues is that we believe that this algorithm might be of interest in the context of implementing it on the simulators of Rabin automata, such as Rabinizer [10].

### v3

@@ 5. Decidability of MSO over the real order with Borel Set Quantifiers
Recall Notation 4.3. We shall translate $\varphi$ into an MSO-sentence $\varphi^{\ast}$ of ${\mathfrak{M}}_{1+\alpha}$. Let $H_{1},\ldots H_{n}$ be the set variables appearing in $\varphi$. We identify individual variables $\rho$ in $\varphi$ with sets $\{\rho\}$, which are non-empty closed sets in $[T]$. In particular they also are $\Sigma_{\alpha}$-sets. Let $\rho_{1},\ldots\rho_{m}$ be all the individual variables appearing in $\varphi$. By applying Lemma 4.8 finitely many times if necessary, we can assume that we have chosen the codes $c(H_{i})$ and $c({\rho}_{j})$ for $i<n,j<m$ within $T^{1+\alpha}]$ so that
@@
Suppose that $\langle F_{\alpha}:\,\alpha<\omega_{1}\rangle$ is a decidable sequence of decidable sentences, namely, there is a recursive procedure which on an input $\alpha$ outputs the truth value of $F_{\alpha}$. Then the statement $(\forall\alpha<\omega_{1})F_{\alpha}]$ is decidable.
@@
As the set of decidable sentences is closed under $\neg,\wedge$ and $\vee$, the nontrivial case is the quantifier case. Suppose first that $\varphi$ has the prenex normal form $(\forall{\mathbf{A}})\psi({\mathbf{A}})$ where $\psi$ does not have any set quantifiers. For each $\alpha<\omega_{1}$ consider the sentence $F_{\alpha}\equiv(\forall{\mathbf{A}})[(\Sigma_{\alpha}({\mathbf{A}})\vee\Pi_{\alpha}({\mathbf{A}})\implies\psi({\mathbf{A}})]$. This is a sentence in the MSO theory of $\mathfrak{C}$ and so decidable by Theorem 5.1. Moreover, considering the sequence $\langle F_{\alpha}:\,\alpha<\omega_{1}\rangle$, it is what we called a decidable sequence of decidable sentences in Theorem 5.2.
@@ 6. Effective Borel Determinacy
A winning strategy $S$ for I is a function defined on $\bigcup_{n<\omega}{}^{2n}2$ such that for any play $\rho$ of $G(H)$ such that at every step $n$ the player I plays $S(\rho\upharpoonright 2n)$, the player I wins (this is a recursive definition). We similarly define a winning strategy for II. The set $H$ is determined if one of the players has a winning strategy in $G(H)$.
@@
Philip Wolfe [21] proved the determinacy of $F_{\sigma}$-sets. In [19, §2.5], Rabin showed that the statement that every $F_{\sigma}$-set is determined, is expressible as an MSO-sentence over ${\mathfrak{N}}_{2}$. He concluded that therefore, Wolfe’s theorem is among those decided by the decision procedure for S2S. In particular, knowing that all $F_{\sigma}$-sets are determined, we can conclude that for every $F_{\sigma}$-set there is a recursive procedure to decide which player wins the game. This recursive procedure is given by a run of the Rabin automaton. This is what we may call an effective version of the determinacy for $F_{\sigma}$-sets.
@@ 7. Closing Remarks
The elegant proof of the Borel Determinacy given by Kechris in [9, Th, 20.5] relates to Martin’s paper [15]. Martin mentions in [15] that it was proposed by Yannis Moschovakis in his book [18] that such a proof should exist, although it was not found then. Before this proof, some earlier papers on the subject used a different technique. The idea was to associate to each Borel game $G$ a game $G^{\ast}$ on an open set in a different Polish space, so that if the player I or II has a winning strategy in $G^{\ast}$, then the player I or II respectively has a winning strategy in $G$. The Borel Determinacy then follows from the Gale-Stewart theorem on the determinacy of Determinacy Gales on open sets in Polish spaces [6] (this would require proper formulation of such games, more general than Definition 6.1). The idea of translation between $G$ and $G^{*}$ was introduced by Martin in [13], where he claims the proof of Borel Determinacy from the existence of an $\omega_{1}$-Erdős cardinal. His paper [14] claims Borel determinacy in ZFC, using a better coding. In the two papers the details of the proofs are missing or incorrect. With this in mind, we have tried to prove Borel Determinacy in yet another way, to match Martin’s original intuition. This would be by playing games on codes for $\Sigma_{\alpha}\cup\Pi_{\alpha}$-sets, understanding those as closed games in an appropriate Polish subspace of $T^{1+\alpha}]$ (see §2.1) and then projecting to the ordinary games on closed sets. The intended proof is by induction on $\alpha$ and uses a direct proof for $F_{\sigma}$-determinacy given in [9, Exercise 20.9] at the non-trivial part of the projection between games in the successor case. This proof would correspond to the original programme of proving Borel Determinacy as intended by Martin in his papers [13] and [14]. We have not worked out all of the details of this proof, as from the purely mathematical point of view it seems unnecessary to give a very long combinatorial argument for a theorem whose proof is already known and well written ([9, Th 20.5]). In fact, we believe that the combinatorial proof is simply a recounting of the decision procedure for the player which has the winning strategy. The part of the combinatorial proof that we did write is available in the code of this paper on the ArXiv. If an interested person uses it, they should ask for a permission and their work should have a proper mention of the source (this paper). The reason we mention these issues is that we believe that this algorithm might be of interest in the context of implementing it on the simulators of Rabin automata, such as Rabinizer [10].